OpenStudy (anonymous):

What is the relationship between the lines determined by these two equations? 3x + 6y = 8 y = 2x – 8

OpenStudy (welshfella):

rearrange the first equation so its in the form y = mx + c

OpenStudy (welshfella):

can you do that?

OpenStudy (anonymous):

parallel perpendicular neither parallel nor perpendicular they are the same line are the options lol

OpenStudy (campbell_st):

the 2nd line is in slope intecept form can you take the 1st line and make y the subject..?

OpenStudy (welshfella):

if you do the rearranging you'll see which one it is

OpenStudy (anonymous):

whaaaaat..

OpenStudy (anonymous):

Idgi..

OpenStudy (welshfella):

yes y = mx + c is slope intercept form

OpenStudy (campbell_st):

ok... 3x + 6y = 8 -3x -3x subtract 3x from both sides of the equation --------------------- 6y = -3x + 8 how do you get a single y on the left side of the equation..?

OpenStudy (welshfella):

I've gotta go I'm sure campbell can help you

OpenStudy (campbell_st):

you will need to divide by a value in the next step

OpenStudy (anonymous):

okay..thanks anyway (:

OpenStudy (anonymous):

it whatever campbell lol I'll just guess. thanks anyway.